Solve the following equations.  14y – 8 = 13
Solve the following equations.  14y – 8 = 13

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

14y–8=13
14y = 13+8
14y=21
y = 21/14
y=3/2

Thanks